# The Holomorphic Extension Property for $k$-du Bois Singularities

### v2

@@ Abstract.
Let $X$ be a normal complex algebraic variety. We show that the holomorphic extension property holds in degree $p<\mathrm{codim}_{X}(X_{\mathrm{sing}})$ when $X$ has du Bois singularities, giving an improvement on Flenner’s criterion for arbitrary singularities. As an application, we study the $k$-du Bois definition from the perspective of holomorphic extension and compare how different restrictions on $\mathscr{H}^{0}(\underline{\Omega}_{X}^{p})$ affect the singularities of $X$, where $\underline{\Omega}_{X}^{p}$ is the $p^{th}$-graded piece of the du Bois complex.
@@ 1.1. Holomorphic and Logarithmic Extension
Holomorphic extension is a weak condition for small $p$: if $\Sigma$ is the singular locus of $X$, then (1.1) is an isomorphism for every $10\leq p<\mathrm{codim}_{X}(\Sigma)-$ 1 by Flenner’s criterion [Fle88, Theorem]. Moreover, if holomorphic extension holds in degree $p$, then holomorphic extension holds in degree $k$ for every $0\leq k\leq p$ [KS21, Theorem 1.4]. For $p\geq\mathrm{codim}_{X}(\Sigma)-1$, extension is less topological and (1.1) can be strict: by definition, strictly log-canonical singularities fail holomorphic extension in degree $\dim X$. Instead, it is better to consider a variant of the inclusion (1.1) which allows for logarithmic poles. We say that a normal variety $X$ satisfies the logarithmic extension property in degree $p$ if the inclusion
@@ 1.2. Extension for du Bois Singularities
is a quasi-isomorphism, where $\underline{\Omega}_{X}^{0}:=\mathrm{gr}_{F}^{0}\underline{\Omega}_{X}^{\bullet}$ is the $0^{th}$-graded piece of the du Bois complex $(\underline{\Omega}_{X}^{\bullet},F)$, an object in the category of filtered complexes of $\mathscr{O}_{X}$-modules generalizing the holomorphic de Rham complex of smooth algebraic varieties.
du Bois singularities are an important in the Hodge theory of singularities, as they arise in the construction of Deligne’s mixed Hodge structure [Del74] on the cohomology of algebraic varieties. They also play a fundamental role in the minimal model program, as log-canonical and rational singularities are du Bois, see [KK10, Theorem 1.4] [Kov99, Corollary 2.6], [Sai00, 5.4. Theorem]. We consider then the following questions:
@@
If $X$ is du Bois and Cohen-Macaulay, the answer to Question (ii) is already known: if $\pi:\widetilde{X}\to X$ is a log-resolution of singularities with exceptional divisor $E$, then the sheaf $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive by [KSS10, Theorem 1.1] [^1] In fact, a Cohen-Macaulay variety is du Bois if and only if $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive, giving a natural generalization of Kempf’s criterion for rational singularities., and so $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$ is reflexive for all $0\leq p\leq\dim X$ [KS21, Theorem 1.5]. We can extend the ideas of [KSS10] to give an analytic proof of logarithmic extension for normal du Bois singularities, giving a positive answer to Question (ii):
@@
If $X$ is a normal complex algebraic variety with at worst du Bois singularities, then logarithmic extension holds in all degrees $0\leq p\leq\dim X$: for any logarithmic resolution of singularities $\pi:\widetilde{X}\to X$, the inclusion $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)\hookrightarrow\Omega_{X}^{[p]}$ is an isomorphism.
@@
is an isomorphism for $p=1$, where $\pi:\widetilde{X}\to X$ is a log-resolution of singularities, and showing $\pi_{*}\Omega_{\widetilde{X}}^{1}(\log E)$ is reflexive [GK14]. By Flenner’s criterion, this is novel exactly when $\mathrm{codim}_{X}(\Sigma)=2$. Theorem 1.2 can also be seen to hold as a corollary of Theorem 1.1 for isolated singularities by an old result of van Straten-Steenbrink: if $j:U\hookrightarrow X$ is the inclusion of the regular locus $U$, there is a differentiation morphism
@@
Recalling that general hyperplane sections preserve the du Bois property, an inductive hypothesis would imply that the left and right vertical morphisms are isomorphisms — this is exactly the approach used in [GK14] for $p=1$, but they need to use the negativity lemma [GKK10, Proposition 7.5], which seems particularly special for $p=1$. Induction seems insufficient in proving $\mathscr{O}_{H}\otimes\pi_{*}\Omega_{\widetilde{X}}^{p}\to\mathscr{O}_{H}\otimes\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$ is an isomorphism, and the reflexivity of the sheaf $(\pi|_{H})_{*}\Omega_{\widetilde{H}}^{p}(\log E|_{H})$ is not sufficient to prove the reflexivity of $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$. We require additional input.
@@ 1.4. Extension Criterion and Hodge Modules
We will often apply this to a singular variety $X$ of dimension $n$ by considering a local embedding $X\subset Y$ into a smooth complex manifold $Y$.
To highlight the idea, let $X$ be a normal variety of dimension $n$ with at worst isolated du Bois singularities, which implies that the cohomology sheaves $R^{j}\pi_{*}\mathscr{O}_{\widetilde{X}}(-E)=0$ for $i>0$ for any log-resolution of singularities $\pi:\widetilde{X}\to X$. By Grauert-Rimenschneider vanishing, there is an induced quasi-isomorphism
(1.6) $\pi_{*}\omega_{\widetilde{X}}(E)\cong\mathbf{R}\pi_{*}\omega_{\widetilde{X}}(E),$
and Grothendieck duality immediately implies that
$R^{j}\mathscr{H}om_{\mathscr{O}_{X}}(\pi_{*}\omega_{\widetilde{X}}(E),\omega_{X}^{\bullet})[-n]\cong R^{j}\pi_{*}\mathscr{O}_{\widetilde{X}}(-E)=0$
for $j>0$. By Proposition 1.3, $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive. We emphasize that this vanishing is much stronger than logarithmic extension and is only equivalent if $X$ is Cohen-Macaulay. Theorem 1.1 follows more generally from the following lemma:
Lemma 1.4.
For holomorphic extension, the problem is more subtle. We need to extend Proposition 1.3 to the level of complexes:
@@
The naive approach is to consider the support of the sheaves $R^{j}\pi_{*}\Omega_{\widetilde{X}}^{\dim X-p}$ for $p=\mathrm{codim}_{X}(\Sigma)$, but this seems hopeless: for instance, $R^{1}\pi_{*}\Omega_{\widetilde{X}}^{1}$ is non-zero for ADE surface singularities — these singularities satisfy holomorphic extension in all degrees, either classically or by [KS21, Corollary 1.8]. Instead, we use the brilliant approach of Kebekus-Schnell and the theory of Hodge modules.
@@ 1.5. Log Forms to Holomorphic Forms
1.6. Holomorphic Extension for $k$-du Bois Singularities
du Bois singularities are of fundamental interest to algebraic geometers as a general class of singularities for which deformation theory and Hodge theory can be studied. It is interesting to consider $k$-du Bois singularites, the class of singularities which restricts the higher graded pieces of the du Bois complex
We say that a normal complex algebraic variety $X$ has weakly (or pre-) $k$-du Bois singularities if the cohomology sheaves $\mathscr{H}^{j}\underline{\Omega}_{X}^{p}=0$ for $j>0$ and $0\leq p\leq k$, where $\underline{\Omega}_{X}^{p}:=\mathrm{gr}_{F}^{p}\underline{\Omega}_{X}^{\bullet}[p]$. Alternatively, $X$ has weakly $k$-du Bois singularities if the natural map
@@ 1.6. Holomorphic Extension for $k$-du Bois Singularities
is a quasi-isomorphism. This notion has been studied in [SVV23] and is a generalization of the notion of $k$-du Bois singularities appearing in [MOPW21], [JKSY21], and [MP22].
An interesting question is then to what extent Theorem 1.2 can be improved for weakly $k$-du Bois singularities. For hypersurfaces, the answer is known: a variety $X$ with hypersurface singularities satisfies holomorphic extension in degree $\dim X$ if and only if $X$ is 1-du Bois, which means that $X$ is du Bois and the natural map
@@
is a quasi-isomorphism, where $\Omega_{X}^{1}$ is the sheaf of Kähler 1-forms. This seems to be rather special for hypersurface (or more generally complete intersection) singularities, as we can write down a Cohen-Macaulay variety with weakly $k$-du Bois singularities and $k>0$ in any dimension for which Theorem 1.2 is still optimal (see Example 6.1). The algebraic properties of the sheaf $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ appear to be independent of the vanishing — or non-vanishing — of the sheaves $\mathscr{H}^{j}\underline{\Omega}_{X}^{p}$. Instead, we consider what happens when $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is already reflexive. One result in this direction is the following:
@@
Let $X$ be a normal variety with weakly $k$-du Bois singularities, and let $\pi:\widetilde{X}\to X$ be a resolution of singularities. If $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ is reflexive, then $\pi_{*}\Omega_{\widetilde{X}}^{k+1}$ is reflexive.
@@ 1.7. Higher Rational Singularities and Reflexivity of $\Omega_{X}^{p}$
Let $X$ be a normal complex algebraic variety. Following [MOPW21] and [JKSY21], we say that $X$ has $k$-du Bois singularities if the natural map $\Omega_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $p$, where $\Omega_{X}^{p}$ is the sheaf of Kähler $p$-forms. As a generalization of du Bois singularities, this is very natural, as it comes from a morphism of complexes $\Omega_{X}^{\bullet}\to\underline{\Omega}_{X}^{\bullet}$. However, this definition is rather restrictive: in the case of hypersurface singularities, this forces $\Omega_{X}^{p}$ to be reflexive for every $0\leq p\leq k$ and restricts the codimension of the singularities.
Instead, one can consider what happens when we restrict the duals of the higher du Bois complexes. We say that has weakly $k$-rational singularities if
@@
for each $0\leq p\leq k$, where $\mathbb{D}_{X}$ is the Grothendieck duality functor. Note that there is a quasi-isomorphism $\mathbb{D}_{X}(\underline{\Omega}_{X}^{n})\cong\mathbf{R}\pi_{*}\mathscr{O}_{\widetilde{X}}$, whence 00-rational is the same as rational singularities for normal complex varieties. This is a generalization of the $k$-rational definition established for complete intersections, which requires the natural morphism $\Omega_{X}^{p}\xrightarrow{\sim}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ to be an isomorphism for $0\leq p\leq k$.
As was observed in [MOPW21], the $k$-du Bois definition restricts the codimension of the singularities of a hypersurface for large $k$, and $X$ is smooth if and only if $k>\frac{1}{2}\dim X$. We generalize this result for $k$-rational singularities. We remark that this appears to be the correct generalization, as $k$-du Bois singularities are $(k-1)$-rational for hypersurfaces by [FL22b].
Theorem 1.9.
If $X$ is a normal complex algebraic variety with $k$-rational and Gorenstein singularities, then $k\leq\frac{1}{2}\mathrm{codim}_{X}(\Sigma)$.
1.8. Hodge Theory of Weakly $k$-Rational Singularities
@@
If $X$ is weakly $k$-du Bois, we get a decomposition
$H^{k}(X,\mathbb{C})=\bigoplus_{p+q=k}H^{q}(X,\mathscr{H}^{0}\underline{\Omega}_{X}^{p}).$
In general, restricting the du Bois complex does not affect the weight filtration on the cohomology, as $\underline{\Omega}_{X}^{\bullet}$ does not usually admit the structure of a Hodge module. To get pure Hodge modules, we consider again $k$-rational definition.
@@
If $X$ is a normal complex algebraic variety with weakly $k$-rational singularities, then the Hodge filtration on $H^{k}(X,\mathbb{Q})$ induces a pure Hodge structure.

### v3

# The Holomorphic Extension Property for Higher Du Bois Singularities

@@ Abstract.
Let $X$ be a normal complex algebraic variety. We show that the holomorphic extension property holds in degree $p<\mathrm{codim}_{X}(X_{\mathrm{sing}})$ when $X$ has Du Bois singularities, giving an improvement on Flenner’s criterion for arbitrary singularities. As an application, we study the $m$-Du Bois definition from the perspective of holomorphic extension and compare how different restrictions on $\mathscr{H}^{0}(\underline{\Omega}_{X}^{p})$ affect the singularities of $X$, where $\underline{\Omega}_{X}^{p}$ is the $p^{th}$-graded piece of the Du Bois complex.
@@ 1.1. Holomorphic and Logarithmic Extension
Holomorphic extension is a weak condition for small $p$: if $\Sigma$ is the singular locus of $X$, then (1.1) is an isomorphism for every $10\leq p<\mathrm{codim}_{X}(\Sigma)-$ 1 by Flenner’s criterion [Fle88, Theorem]. Moreover, if holomorphic extension holds in degree $p$, then holomorphic extension holds in degree $k$ for every $0\leq k\leq p$ [KS21, Theorem 1.4]. For $p\geq\mathrm{codim}_{X}(\Sigma)-1$, extension is less topological and (1.1) can be strict: by definition, Gorenstein log-canonical singularities fail holomorphic extension in degree $\dim X$. Instead, it is better to consider a variant of the inclusion (1.1) which allows for logarithmic poles. We say that a normal variety $X$ satisfies the logarithmic extension property in degree $p$ if the inclusion
@@ 1.2. Extension for Du Bois Singularities
is a quasi-isomorphism, where $\underline{\Omega}_{X}^{0}:=\mathrm{gr}^{F}_{0}\underline{\Omega}_{X}^{\bullet}$ is the $0^{th}$-graded piece of the Du Bois complex $(\underline{\Omega}_{X}^{\bullet},F)$, an object in the category of filtered complexes of $\mathscr{O}_{X}$-modules generalizing the holomorphic de Rham complex of smooth algebraic varieties.
Du Bois singularities are an important class of singularities in the Hodge theory of singularities, as they arise in the construction of Deligne’s mixed Hodge structure [Del74] on the cohomology of algebraic varieties. They also play a fundamental role in the minimal model program, as log-canonical and rational singularities are Du Bois, see [KK10, Theorem 1.4] [Kov99, Corollary 2.6], [Sai00, 5.4. Theorem]. We consider then the following questions:
@@
If $X$ is Du Bois and Cohen-Macaulay, the answer to Question (ii) is already known: if $\pi:\widetilde{X}\to X$ is a log-resolution of singularities with exceptional divisor $E$, then the sheaf $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive by [KSS10, Theorem 1.1] [^1] In fact, a Cohen-Macaulay variety is Du Bois if and only if $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive, giving a natural generalization of Kempf’s criterion for rational singularities., and so $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$ is reflexive for all $0\leq p\leq\dim X$ [KS21, Theorem 1.5]. This gives a short proof of a result of Kovács-Graf.
@@
[GK14b, Theorem 4.1] If $X$ is a normal complex algebraic variety with at worst Du Bois singularities, then logarithmic extension holds in all degrees $0\leq p\leq\dim X$: for any logarithmic resolution of singularities $\pi:\widetilde{X}\to X$, the inclusion $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)\hookrightarrow\Omega_{X}^{[p]}$ is an isomorphism.
@@
is an isomorphism for $p=1$, where $\pi:\widetilde{X}\to X$ is a log-resolution of singularities, and showing $\pi_{*}\Omega_{\widetilde{X}}^{1}(\log E)$ is reflexive [GK14a]. By Flenner’s criterion, this is novel exactly when $\mathrm{codim}_{X}(\Sigma)=2$. Theorem 1.2 can also be seen to hold as a corollary of Theorem 1.1 for isolated singularities by an old result of van Straten-Steenbrink: if $j:U\hookrightarrow X$ is the inclusion of the regular locus $U$, there is a differentiation morphism
@@
Recalling that general hyperplane sections preserve the Du Bois property, an inductive hypothesis would imply that the left and right vertical morphisms are isomorphisms — this is exactly the approach used in [GK14a] for $p=1$, but they need to use the negativity lemma [GKK10, Proposition 7.5], which seems particularly special for $p=1$. Induction seems insufficient in proving $\mathscr{O}_{H}\otimes\pi_{*}\Omega_{\widetilde{X}}^{p}\to\mathscr{O}_{H}\otimes\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$ is an isomorphism, and the reflexivity of the sheaf $(\pi|_{H})_{*}\Omega_{\widetilde{H}}^{p}(\log E|_{H})$ is not sufficient to prove the reflexivity of $\pi_{*}\Omega_{\widetilde{X}}^{p}(\log E)$. We require additional input.
@@ 1.4. Extension Criterion and Hodge Modules
We will often apply this to a singular variety $X$ of dimension $n$ by considering a local embedding $X\subset Y$ into a smooth complex manifold $Y$. There is also a derived version of this criterion discussed in [KS21]. We remark that this statement is stronger than holomorphic extension:
@@
To highlight the idea, let $X$ be a normal variety of dimension $n$ with at worst isolated Du Bois singularities, which implies that the cohomology sheaves $R^{j}\pi_{*}\mathscr{O}_{\widetilde{X}}(-E)=0$ for $i>0$ for any log-resolution of singularities $\pi:\widetilde{X}\to X$. We use Proposition 1.4 to test holomorphic extension on
(1.6) $\pi_{*}\omega_{\widetilde{X}}(E)=R^{0}\pi_{*}\omega_{\widetilde{X}}(E).$
By Grothendieck duality,
$R^{j}\mathscr{H}om_{\mathscr{O}_{X}}(\mathbf{R}\pi_{*}\omega_{\widetilde{X}}(E),\omega_{X}^{\bullet})[-n]\cong R^{j}\pi_{*}\mathscr{O}_{\widetilde{X}}(-E)=0$
for $j>0$. Thus, $\pi_{*}\omega_{\widetilde{X}}(E)$ is reflexive. We emphasize that this vanishing is much stronger than logarithmic extension and is only equivalent if $X$ is Cohen-Macaulay. Theorem 1.1 follows more generally from the following lemma:
Lemma 1.5.
For holomorphic extension, the problem is more subtle. The naive approach is to consider the support of the sheaves $R^{j}\pi_{*}\Omega_{\widetilde{X}}^{\dim X-p}$ for $p=\mathrm{codim}_{X}(\Sigma)$, but this seems hopeless: for instance, $R^{1}\pi_{*}\Omega_{\widetilde{X}}^{1}$ is non-zero for ADE surface singularities — these singularities satisfy holomorphic extension in all degrees, either classically or by [KS21, Corollary 1.8]. Instead, we use the brilliant approach of Kebekus-Schnell and the theory of Hodge modules.
@@ 1.5. Log Forms to Holomorphic Forms
1.6. Holomorphic Extension for $m$-Du Bois Singularities
Du Bois singularities are of fundamental interest to algebraic geometers as a general class of singularities for which deformation theory and Hodge theory can be studied. It is interesting to consider $m$-Du Bois singularities, the class of singularities which restricts the higher graded pieces of the Du Bois complex.
We say that a normal complex algebraic variety $X$ has weakly (or pre-) $m$-Du Bois singularities if the cohomology sheaves $\mathscr{H}^{j}\underline{\Omega}_{X}^{p}=0$ for $j>0$ and $0\leq p\leq m$, where $\underline{\Omega}_{X}^{p}:=\mathrm{gr}^{F}_{p}\underline{\Omega}_{X}^{\bullet}[p]$. Alternatively, $X$ has weakly $m$-Du Bois singularities if the natural map
@@ 1.6. Holomorphic Extension for $m$-Du Bois Singularities
is a quasi-isomorphism. This notion has been studied in [SVV23] and is a generalization of the notion of $m$-Du Bois singularities appearing in [MOPW21], [JKSY21], and [MP22].
An interesting question is then to what extent Theorem 1.2 can be improved for weakly $m$-Du Bois singularities. For hypersurfaces, the answer is known: a variety $X$ with hypersurface singularities satisfies holomorphic extension in degree $\dim X$ if $X$ is 1-Du Bois, which means that $X$ is Du Bois and the natural map
@@
is a quasi-isomorphism, where $\Omega_{X}^{1}$ is the sheaf of Kähler 1-forms. This seems to be rather special for hypersurface (or more generally complete intersection) singularities, as we can write down a Cohen-Macaulay variety with weakly $m$-Du Bois singularities and $k>0$ in any dimension for which Theorem 1.2 is still optimal (see Example 6.1). The algebraic properties of the sheaf $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ appear to be independent of the vanishing — or non-vanishing — of the sheaves $\mathscr{H}^{j}\underline{\Omega}_{X}^{p}$. Instead, we consider what happens when $\mathscr{H}^{0}\underline{\Omega}_{X}^{p}$ is already reflexive. One result in this direction is the following:
@@
Let $X$ be a normal variety with weakly $m$-Du Bois singularities, and let $\pi:\widetilde{X}\to X$ be a resolution of singularities. If $\mathscr{H}^{0}\underline{\Omega}_{X}^{k}$ is reflexive, then $\pi_{*}\Omega_{\widetilde{X}}^{k+1}$ is reflexive.
@@ 1.7. Higher Rational Singularities and Reflexivity of $\Omega_{X}^{p}$
Let $X$ be a normal complex algebraic variety. Following [MOPW21] and [JKSY21], we say that $X$ has $m$-Du Bois singularities if the natural map $\Omega_{X}^{p}\to\underline{\Omega}_{X}^{p}$ is a quasi-isomorphism for each $p$, where $\Omega_{X}^{p}$ is the sheaf of Kähler $p$-forms. As a generalization of Du Bois singularities, this is very natural, as it comes from a morphism of complexes $\Omega_{X}^{\bullet}\to\underline{\Omega}_{X}^{\bullet}$. However, this definition is rather restrictive: in the case of hypersurface singularities, this forces $\Omega_{X}^{p}$ to be reflexive for every $0\leq p\leq k$ and restricts the codimension of the singularities.
Instead, one can consider what happens when we restrict the duals of the higher Du Bois complexes. We say that has weakly $m$-rational singularities if
@@
for each $0\leq p\leq k$, where $\mathbb{D}_{X}$ is the Grothendieck duality functor. Note that there is a quasi-isomorphism $\mathbb{D}_{X}(\underline{\Omega}_{X}^{n})\cong\mathbf{R}\pi_{*}\mathscr{O}_{\widetilde{X}}$, whence 00-rational is the same as rational singularities for normal complex varieties. This is a generalization of the $m$-rational definition established for complete intersections, which requires the natural morphism $\Omega_{X}^{p}\xrightarrow{\sim}\mathbb{D}_{X}(\underline{\Omega}_{X}^{n-p})$ to be an isomorphism for $0\leq p\leq k$.
The $m$-rational property is very strong, as it eventually implies $\Omega_{X}^{1}$ is maximal Cohen-Macaulay. The MCM property has been extensively studied in the literature, and the MCM property for the module of Kähler differentials has been looked at in the case of hypersurfaces and when $\mathrm{pd}~{}\Omega_{X,x}^{p}<\infty$. We give one new result using the theory of Hodge modules.
Proposition 1.9.
If $X$ is an $n$-fold Gorenstein variety with at worst quotient singularities, then $\Omega_{X}^{p}$ is reflexive for some $1\leq p\leq n$ if and only if $X$ is smooth.
1.8. Hodge Theory of Weakly $m$-rational Singularities
@@
If $X$ is weakly $m$-Du Bois, we get a decomposition
$H^{k}(X,\mathbb{C})\cong\bigoplus_{p+q=k}H^{q}(X,\mathscr{H}^{0}\underline{\Omega}_{X}^{p}).$
In general, restricting the Du Bois complex does not affect the weight filtration on the cohomology, as $\underline{\Omega}_{X}^{\bullet}$ does not usually admit the structure of a Hodge module. To get pure Hodge modules, we consider again $m$-rational definition.
@@
If $X$ is a normal and proper complex algebraic variety with weakly $m$-rational singularities, then the Hodge filtration on $H^{k}(X,\mathbb{Q})$ induces a pure Hodge structure.